# On Generating Monolithic and Model Reconciling Explanations in Probabilistic Scenarios

### v1

@@ 7 Concluding Remarks
In this paper, we started by describing a framework for generating probabilistic monolithic explanations within uncertain knowledge bases (e.g., belief bases). Our work is an attempt to bridge the gap between classical explanation models and the inherent uncertainty that characterize real-world scenarios. We introduced the concepts of explanatory gain and explanatory power as quantitative measures to evaluate the effectiveness and relevance of explanations, thereby offering a better understanding of explanation quality.
Additionally, we presented an extension to the model reconciliation problem for generating probabilistic model reconciling explanations, which addresses the crucial need for reconciling model differences between an agent and a human model, specifically when the human model is not known with certainty. While we considered a classical knowledge base as the model of the agent in this work, future work could look into the assumption that the agent model is also probabilistic (i.e., a belief base). Reconciling two belief bases poses a significant challenge due to the complexity inherent in probabilistic logic and the need to account for uncertainty in both models. This task involves aligning two sets of beliefs, each associated with its own probability distributions, to achieve a coherent understanding that reflects the true state of affairs or intentions accurately. The process must consider not only the logical consistency between the belief bases but also the probabilistic coherence, ensuring that the reconciled belief base maintains a meaningful probability distribution over its assertions.
Furthermore, we developed algorithms that leverage the duality between minimal correction sets (MCSes) and minimal unsatisfiable subsets (MUSes) and demonstrated their potential for generating probabilistic explanations. Our experimental evaluations across diverse benchmarks underscore the practicality and effectiveness of these algorithms, validating their applicability in real-world settings. While our algorithms are presented within the framework of propositional logic, their underlying principles are broadly applicable to any constraint system where the satisfiability of subsets can be decided. This versatility opens avenues for extending our work to other logical systems, such as Markov Logic Networks (?) and Probabilistic Logic Programs (?).

### v2

@@ 10 Concluding Remarks
In this paper, we attempted to bridge the gap between classical explanation models and the inherent uncertainty found in real-world scenarios. We started by describing a framework for generating probabilistic monolithic explanations within uncertain knowledge bases (e.g., belief bases), and introduced the concepts of explanatory gain and explanatory power as quantitative measures to evaluate the effectiveness and relevance of explanations, thus offering a better characterization of explanation quality. Additionally, we presented an extension to the model reconciliation problem for generating probabilistic model reconciling explanations, which addresses the need for reconciling model differences between an agent and a human model, specifically when the human model is not known with certainty.
Furthermore, we developed algorithms that leverage the duality between minimal correction sets (MCSes) and minimal unsatisfiable subsets (MUSes) and demonstrated their potential for generating probabilistic explanations. Our experimental evaluations across different benchmarks underscore the effectiveness of these algorithms, suggesting they are promising in practical scenarios.
Acknowledgments
We thank the anonymous reviewers, whose suggestions improved the quality of our paper. Stylianos Loukas Vasileiou and William Yeoh are partially supported by the National Science Foundation (NSF) under award 2232055. Son Tran is partially supported by NSF grants 1914635 and 2151254 as well as ExpandAI grant 2025-67022-44266. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the United States government.
A Illustrative Example: Simple Package Delivery Problem
In Section 4, we presented a simple probabilistic planning scenario (?) that can be used by our framework. This problem can be fully specified in the form of the probabilistic planning domain definition language (PPDDL) (?), as shown in Listings 1 and 2. Given the PPDDL descriptions, we can then encode this problem into (propositional) probabilistic logic (?), similarly to the encoding for classical planning problems (cf. ? (?)). Specifically, for $l,l^{\prime}\in\{room1,room2\}$ and $c\in\{A,B\}$, we can encode the belief base:
Initial States: (Starting states and belief about the environment)
$(\texttt{robot}\textrm{-}\texttt{at}(room1)_{0},\infty)$
$(\neg\texttt{robot}\textrm{-}\texttt{at}(room2)_{0},\infty)$
$(\neg\texttt{package}\textrm{-}\texttt{delivered}_{0},\infty)$
$(\texttt{crowded}(A),w_{1})$
$(\texttt{crowded}(B),w_{2})$
Goal State: (Final state to be reached)
$(\texttt{package}\textrm{-}\texttt{delivered}_{t_{n}},\infty)$
Action Preconditions: (Fluents that must be true at timestep $t$ for the action to be executed)
$(\texttt{move}(l,l^{\prime},c)_{t}\rightarrow\texttt{robot}\textrm{-}\texttt{at}(l)_{t},\infty)$
$(\texttt{deliver}(l)_{t}\rightarrow\texttt{robot}\textrm{-}\texttt{at}(l)_{t},\infty)$
Deterministic Action Effects: (Fluents that become true at timestep $t+1$ if the action executed at timestep $t$)
$(\texttt{deliver}(l)_{t}\rightarrow\texttt{package}\textrm{-}\texttt{delivered}_{t+1},\infty)$
Probabilistic Action Effects: (Fluents that become true at timestep $t+1$ with a certain probability if the action executed at timestep $t$)
$(\texttt{move}(l,l^{\prime},c)_{t}\wedge\texttt{crowded}(c)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(l^{\prime})_{t+1}\wedge\lnot\texttt{robot}\textrm{-}\texttt{at}(l)_{t+1},w_{3})$
$(\texttt{move}(l,l^{\prime},c)_{t}\wedge\texttt{crowded}(c)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(l)_{t+1},w_{4})$
$(\texttt{move}(l,l^{\prime},c)_{t}\wedge\lnot\texttt{crowded}(c)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(l^{\prime})_{t+1}\wedge\lnot\texttt{robot}\textrm{-}\texttt{at}(l)_{t+1},w_{5})$
$(\texttt{move}(l,l^{\prime},c)_{t}\wedge\lnot\texttt{crowded}(c)\rightarrow\texttt{robot}\textrm{-}\texttt{at}(l)_{t+1},w_{6})$
Explanatory Frame Axioms: (Fluents do not change between subsequent timesteps $t$ and $t+1$ unless they are effects of actions that are executed at timestep $t$)
$(\texttt{robot}\textrm{-}\texttt{at}(l)_{t}\wedge\lnot\texttt{robot}\textrm{-}\texttt{at}(l)_{t+1}\rightarrow\texttt{move}(l,l^{\prime},A)_{t}\vee\texttt{move}(room1,room2,B)_{t},\infty)$
$(\lnot\texttt{robot}\textrm{-}\texttt{at}(l)_{t}\wedge\texttt{robot}\textrm{-}\texttt{at}(l)_{t+1}\rightarrow\texttt{move}(l^{\prime},l,A)_{t}\vee\texttt{move}(l^{\prime},l,B)_{t},\infty)$
$(\texttt{package}\textrm{-}\texttt{delivered}_{t}\rightarrow\texttt{package}\textrm{-}\texttt{delivered}_{t+1},\infty)$
$(\lnot\texttt{package}\textrm{-}\texttt{delivered}_{t}\wedge\texttt{package}\textrm{-}\texttt{delivered}_{t+1}\rightarrow\texttt{deliver}(l)_{t},\infty)$
Action exclusions: (Only one action can occur at each timestep $t$)
$(\neg\texttt{move}(l,l^{\prime},A)_{t}\vee\neg\texttt{move}(l,l^{\prime},B)_{t},\infty)$
$(\neg\texttt{move}(l,l^{\prime},c)_{t}\vee\neg\texttt{deliver}(l)_{t},\infty)$
Finally, we can extract a plan by finding an assignment of truth values that satisfies the belief base (i.e., for all timesteps t=0,,n1t=0,...,n-1, there will be exactly one action $a$ such that $a_{t}=True$).
Note that a formula’s weight represents the log odds comparing worlds where the formula is true versus where it is false, assuming all else remains unchanged. However, when formulae share variables with each other, it is impossible to flip one formula’s truth value without affecting others. Despite that, we can still determine appropriate weights through collective formula probabilities. By treating desired probabilities as empirical frequencies for maximum likelihood estimation (MLE) (?), we can derive optimal weights. This approach involves specifying how often each formula should be true, using these as observed frequencies, and applying standard MLE algorithms (e.g., (?)) to compute weights.
Listing 1: PPDDL Domain of Office Robot Delivery.
(define (domain Office-Robot-Delivery)
(:requirements :strips :typing :probabilistic-effects)
(:types
location corridor)
(:predicates
(robot-at ?l - location)
(connected ?l1 ?l2 - location ?c - corridor)
(package-delivered)
(crowded ?c - corridor)
(:action move
:parameters (?from ?to - location ?c - corridor)
:precondition (and
(robot-at ?from)
(connected ?from ?to ?c))
:effect (and
(when (crowded ?c)
(probabilistic
p (and (robot-at ?to) (not (robot-at ?from)))
1-p (and (robot-at ?from))))
(when (not (crowded ?c))
(probabilistic
p (and (robot-at ?to) (not (robot-at ?from)))
1-p (and (robot-at ?from)))))
(:action deliver
:parameters (?l - location)
:precondition (robot-at ?l)
:effect (package-delivered)
Listing 2: PPDDL Problem of Office Robot Delivery.
(define (problem office-delivery)
(:domain Office-Robot-Delivery)
(:objects
room1 room2 - location
A B - corridor)
(:init
(robot-at room1)
(connected room1 room2 A)
(connected room1 room2 B)
(probabilistic p (crowded A))
(probabilistic p (crowded B))
(:goal (package-delivered room2))